OpenStudy (anonymous):

Help please! Evaluate the indefinite integral... x^5((7+9x^6)^11) dx

OpenStudy (anonymous):

u substitution works

OpenStudy (anonymous):

\[u = 7+9x^6 \]du = ?

OpenStudy (cwrw238):

yes - i just looked it up - use substitution

OpenStudy (cwrw238):

can you work out du ?

OpenStudy (anonymous):

so it would be... u=7+9x^6 du=54x^5 (1/54)=x^5

OpenStudy (cwrw238):

if u = ax^n du/dx = an x^(n-1) have you seen this before?

OpenStudy (cwrw238):

du / dx = 54 x^5 dx = du / 54x^5 so we get